OpenStudy (anonymous):

HELP! Find the exact value of cos(11π/12). Hint: begin by rewriting 11π/12 in terms of familiar angles.

OpenStudy (dls):

\[\Huge \cos(\frac{11\pi}{12})=\cos(\pi-\frac{\pi}{12})\]

OpenStudy (anonymous):

cos(x+y)= cos x cosy - sin x sin y use this identity

OpenStudy (anonymous):

11π/12 = 8π/12 + 3π/12 = 2π/3 + π/4 i would use this

OpenStudy (anonymous):

I really don't get how any of that helps me.

OpenStudy (jdoe0001):

\(\bf \huge \frac{11\pi}{12} \implies \frac{8\pi}{12}+\frac{3\pi}{12} \implies \frac{2\pi}{3}+\frac{\pi}{4}\)

OpenStudy (jdoe0001):

http://s4.hubimg.com/u/5401027_f260.jpg

OpenStudy (anonymous):

The answer choices are: √6+√2/4 √3+2/4 √3-2?4 -√ 2-√6/4

OpenStudy (jdoe0001):

so, add them then

OpenStudy (anonymous):

Use the identity I have posted above. cos(x+y)= cos x cosy - sin x sin y

OpenStudy (jdoe0001):

and your Unit Circle

OpenStudy (anonymous):

Oh man, I got to go thanks for the help, I'll solve it when I get back, I still gave you the medal.